Solve the system.
3x+4y=-3
2x+y=8

Answer by funmath(2873)

Solve the system.
E1) 3x+4y=-3
E2) 2x+y=8
This is solving by substitution, since you didn't specify.
Solve E2) for y and substitute that into E1 and solve for x.
E2) -2x+2x+y=-2x+8 -->>y=-2x+8
E1) 3x+4(-2x+8)=-3
3x-8x+32=-3
-5x+32=-3
-5x+32-32=-3-32
-5x=-35
-x/-5=-35/-5
x=7
Substitute that in for x in E2 and solve for y
2(7)+y=8
14+y=8
-14+14+y=8-14
y=-6
The solution is: (x,y)=(7,-6)
Happy Calculating!!!